OpenStudy (openstudygirl2):

Alina has just received 3 video games for her birthday to add to her existing 12 video games. A week later she went to the thrift store and doubled her video games. Another week later she bought 5 more games. Write a numerical expression with parentheses for the total number of video games Alina has now. Do not simplify your expression. Please help will fan and medal

Awolflover1 (awolflover1):

okay so she got 3 new games and added to her existing 12 games and then got 5 more

OpenStudy (openstudygirl2):

20 games

Awolflover1 (awolflover1):

hmm interesting question

Awolflover1 (awolflover1):

\[12(3+5) \] hmm kinda stuck on this one

OpenStudy (openstudygirl2):

....... i have no clue about 12(3+5)

OpenStudy (openstudygirl2):

.............

Awolflover1 (awolflover1):

well it does say ''Write a numerical expression with parentheses for the total number of video games Alina has now.''

OpenStudy (openstudygirl2):

ya and

Awolflover1 (awolflover1):

did you add all the numbers up?

OpenStudy (openstudygirl2):

yes

Awolflover1 (awolflover1):

what did you get?

OpenStudy (openstudygirl2):

but i never doubled 20 so the answer would be 40

OpenStudy (openstudygirl2):

............

Awolflover1 (awolflover1):

40 ?

OpenStudy (openstudygirl2):

yup

Awolflover1 (awolflover1):

are you sure? 3+12+5 = 40 ?

OpenStudy (openstudygirl2):

noooo this is it

Awolflover1 (awolflover1):

oh! i didn't even see that haha good eye

Awolflover1 (awolflover1):

Hmm kinda stuck with this one :\

OpenStudy (raden):

3 + 12 = 15 Doubled each video, it means 15 x 2 =30 And then add 5 video games, 30 + 5 = 35

OpenStudy (robtobey2):

\[2 (12+3)+5=35 \]

OpenStudy (openstudygirl2):

so whats the answer